how- do you solve the following equation:
1/2y = -7

Answer by Alan3354(69443)

how- do you solve the following equation:
1/2y = -7
------------
If you mean
(1/2)y = -7
Multiply by 2.
===============
If you mean
1/(2y) = -7
Multiply by 2y
Divide by -14